# Spin Polarization of $\boldsymbol{\Lambda}$ hyperons from Dissipative Spin Hydrodynamics

### v1

@@ Abstract
We present a framework for spin dynamics in the quark-gluon plasma created in relativistic heavy-ion collisions. Under the approximation of small polarization, macroscopic spin degrees of freedom decouple from the background, and their evolution equations and transport coefficients have been computed using quantum kinetic theory of massive particles with non-local collisions. Employing this theory, we numerically solve dissipative relativistic spin hydrodynamics for the first time. We explore three interaction scenarios between constituent particles and apply this framework to compute both global and local spin polarization of $\Lambda$ hyperons in Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV. Our results show that the initially vanishing spin potential relaxes toward thermal vorticity, driving global polarization. Furthermore, we demonstrate that the sign of longitudinal polarization is sensitive to the interaction type, emphasizing the need for a consistent treatment of dissipative effects in spin hydrodynamics to describe experimental data.
@@
Introduction.– The study of spin polarization in relativistic heavy-ion collisions (HIC) has gained considerable interest in recent years, owing to the observation of global and local polarization of hyperons [1, 2, 3, 4, 5, 6, 7, 8]. Spin polarization observables provide probes of the vorticity and spin dynamics in the hot and dense fireball formed in relativistic HIC. The earliest attempt using a perturbative QCD-inspired model [9] to understand global polarization attributed it to the large orbital angular momentum, which polarizes the quarks in the same direction through spin-orbit coupling, analogous to the well-known Barnett effect [10].
In the quantum-statistical approach, the spin degrees of freedom are assumed to be in local thermodynamic equilibrium at the fireball freezeout, making it possible to define a density operator and derive an analytical expression for spin polarization at leading order in terms of the thermal vorticity, $\varpi^{\mu\nu}=\frac{1}{2}(\partial^{\nu}\beta^{\mu}-\partial^{\mu}\beta^{\nu})$, where $\beta^{\mu}=u^{\mu}/T$ is the inverse four-temperature, with $u^{\mu}$ denoting the fluid four-velocity and $T$ the temperature [11]. The spin polarization generated solely from thermal vorticity shows very good agreement with experimental data on the global polarization of $\Lambda$-hyperons across Beam Energy Scan collision energies [12]. However, the data on local polarization, particularly the longitudinal polarization, exhibits an opposite sign compared to hydrodynamic predictions. It was soon realized that the thermal shear $\xi^{\mu\nu}=\frac{1}{2}(\partial^{\nu}\beta^{\mu}+\partial^{\mu}\beta^{\nu})$ provides an additional contribution to the spin polarization [13, 14, 15, 16, 17]. The combined effects of thermal vorticity and thermal shear successfully reproduce both the magnitude and the correct sign of the local polarization at $\sqrt{s_{NN}}=200$ GeV under the so-called isothermal approximation, where gradients of temperature are neglected on the freezeout surface [14].
@@
While the equations of spin hydrodynamics can be formulated in several ways [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53], the most popular approach is to derive them from quantum kinetic theory (QKT), which describes a quantum-field theoretical system in the limit where the fields are sufficiently localized to be treated in terms of quasi-particles. Utilizing this framework, one possible approach consists in taking the spin tensor to be conserved [26, 27], implying that the energy-momentum tensor is symmetric. A numerical solution of spin hydrodynamics derived in such a way was presented in Ref. [54] with the spin tensor taken to be in de Groot, van Leeuwen, and van Weert (GLW) pseudogauge. For a suitable choice of initial spin potential, their results show that using a delayed initial time for spin hydrodynamics can successfully describe polarization measurements. This is suggestive of strong dissipative effects during the initial stages of the fireball that may be responsible for the relaxation of the spin potential towards its equilibrium value given by the thermal vorticity. Indeed, it is also possible to derive these dissipative effects in a quantum kinetic framework. In Ref. [24], it has been found that the relaxation of the spin potential towards the thermal vorticity is sourced by so-called non-local collisions [55, 56]. Intuitively, this is sensible since, in order to exchange orbital angular momentum and spin, the collision cannot occur at a single spacetime point [27]. On a more technical side, these collisions essentially depend on the spin-changing matrix elements of the underlying theory, and the equations solved in Ref. [54] refer to a system in which these interactions can be neglected. In this work, we go beyond this approximation and, for the first time, numerically solve the equations of dissipative spin hydrodynamics.
@@
Theoretical background.– The degrees of freedom quantifying the spin density of the fluid can be described via the spin potential $\Omega^{\mu\nu}_{0}$, which constitutes an antisymmetric rank-two tensor and can thus be decomposed into electric and magnetic parts, [^1] Note that the sign of $\kappa_{0}^{\mu}$ differs between Ref. [27] and this work, which uses the same conventions as Refs. [24, 53].
@@
Furthermore, when considering dissipative spin hydrodynamics, it has been shown in Ref. [53] that one has to consider an additional rank-2 tensor $\mathfrak{t}^{\mu\nu}$ which has the same symmetries as the shear-stress tensor $\pi^{\mu\nu}$ and can thus be referred to as the spin-shear stress. In the GLW pseudogauge, it appears in the spin tensor as $S^{\mu,\alpha\beta}\sim\mathfrak{t}^{\mu[\alpha}u^{\beta]}$.
@@
The relaxation times $\tau_{\omega}$ and $\tau_{\kappa}$ quantify the time scales on which $\omega_{0}^{\mu}$ and $\kappa_{0}^{\mu}$ approach their asymptotic values, i.e., the first terms on the right-hand sides of Eqs. (2a) and (2b). As mentioned in the introduction, these times are (inversely) proportional to the non-local collision terms. On the other hand, the coefficient $\tau_{\mathfrak{t}}$, which sets the time scale of $\mathfrak{t}^{\mu\nu}$, is (inversely) proportional to local collisions, indicating that the relaxation of the spin-shear stress is not driven solely through spin-orbit coupling. The asymptotic value of the spin-shear stress, given by the first term on the right-hand side of Eq. (2c), is proportional to the shear tensor, reminiscent of the thermal-shear contribution in Ref. [13]. The corresponding transport coefficient $\mathfrak{d}$ is given by a ratio of non-local and local collision terms. In particular, while it depends on the underlying microscopic interaction, it is independent of its absolute strength, and vanishes when the collisions are local.
@@
Figure 1: Time evolution of the energy-weighted space-averaged (left) $xz$- component of the spin potential, and (right) $zz$-component of the spin-shear stress, for three different interactions. Due to the symmetry of our initial condition model, the spatial average of all other components of the spin potential is zero.
@@
Numerical framework.– Solving Eqs. (2a)-(2c) requires $T$, $u^{\mu}$, and their gradients from the background hydrodynamics. We use the code developed in Refs. [54, 58, 59] to solve the equations of background hydrodynamics numerically in (3+1)-dimensions in Milne coordinates $(\tau,x,y,\eta_{s})$, where $\tau=\sqrt{t^{2}-z^{2}}$ and $\eta_{s}=\frac{1}{2}\ln\left(\frac{t+z}{t-z}\right)$. The code, developed in FORTRAN, incorporates the Godunov-type relativistic Harten-Lax-van Leer-Einfeldt (HLLE) approximate Riemann solver for the computation of numerical fluxes at fluid cell boundaries, as described in Refs. [60, 61]. The implementation of Eqs. (2a)-(2c) is inspired by the publicly available vhlle code [61, 62]. Additionally, the FORTRAN code supports parallel programming on multiple threads using OpenMP [63], a feature not present in the public version of vhlle. The equations of background hydrodynamics are closed with an equation of state for which we use the lattice-QCD-based equation of state NEOS-BQS [64, 65, 66, 67, 68, 69]. We extend the code to solve the spin equations, Eqs. (2a)-(2c), with the details provided in Appendix D.
@@
For the computation of the transport coefficients in Eqs. (2a)-(2c), one has to provide an explicit model of the system. Deferring the use of a kinetic framework with all particle species included to future work, we employ a simplified treatment by assuming a system of spin-$\nicefrac{{1}}{{2}}$ particles with a mass of 300 MeV, which is on the order of the thermal masses $m_{\text{th}}\sim gT$ of the particles in the hot and dense fireball. We consider three types of interactions among the particles, namely (i) scalar, where the interaction Lagrangian is $\mathcal{L}_{\text{int,S}}=G(\bar{\psi}\psi)^{2}$, (ii) scalar+pseudoscalar, where $\mathcal{L}_{\text{int,SP}}=G[(\bar{\psi}\psi)^{2}-(\bar{\psi}\gamma_{5}\psi)^{2}]$, [^2] This interaction is equivalent to the well-known Nambu-Jona–Lasinio (NJL) model [70]. and (iii) vector, with $\mathcal{L}_{\text{int,V}}=-G(\bar{\psi}\gamma^{\mu}\psi)^{2}$, with a coupling constant $G$. We denote these three scenarios by $I_{S}$, $I_{SP}$, and $I_{V}$, respectively. As mentioned earlier, the transport coefficient $\mathfrak{d}$ depends only on the form of the interaction but is independent of the coupling constant $G$. The same can be said for the relaxation times $\tau_{\omega}$, $\tau_{\kappa}$, and $\tau_{\mathfrak{t}}$, if we express them in units of the shear-stress relaxation time $\tau_{\pi}$. Consequently, we tabulate the relaxation times $\tau_{\omega}$, $\tau_{\kappa}$, and $\tau_{\mathfrak{t}}$ as functions of $z=m/T$ in units of $\tau_{\pi}$ for the three types of interactions mentioned above (see Fig. 4 in Appendix B). All other transport coefficients in the spin equations are tabulated as functions of $z$ in units of the corresponding spin relaxation times and isotropic pressure to make them dimensionless (see Ref. [53] for details). We assume that these dimensionless ratios are universal, depending only on the nature of the interaction between particles, and thus remain unchanged for the quark-gluon fluid formed in relativistic HIC. Consequently, for transport coefficients that involve pressure, we use the local thermodynamic equilibrium pressure, $P$, of the fireball to extract their numerical values from tables of dimensionless ratios. This approach aligns with standard hydrodynamic simulations, such as MUSIC [71], where second-order transport coefficients computed using kinetic theory in the massless limit are assumed to apply to the quark-gluon fluid. To solve Eqs. (2a)-(2c), we initialize $\omega_{0}^{\mu}$, $\kappa_{0}^{\mu}$, and $\mathfrak{t}^{\mu\nu}$ to zero.
@@
Figure 2: Comparison of results from dissipative spin hydrodynamics with experimental data for three interaction scenarios for Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV. The simulations are performed with fixed impact parameter, $b=8.4$ fm, corresponding to 20-50% centrality. The top panel shows the (left) pseudorapidity and (right) transverse momentum dependence of the $y$-component of $\Lambda$ polarization. The bottom panel shows the (left) azimuthal and (right) transverse momentum dependence of the $z$-component of $\Lambda$ polarization. Experimental data are taken from Refs. [2, 72] with the updated decay parameter ($\alpha_{\Lambda}=0.732$).
@@
The spin vector in the rest frame of $\Lambda$, $S^{*\mu}=(0,\textbf{S}^{*})$ is obtained from $^{\mu}=(S^{0},\textbf{S})$ S ) as follows:
@@
Results.– With the framework described above, we analyze the behavior of dissipative spin hydrodynamics in HIC. In Fig. 1, we show the time evolution of the energy-weighted, space-averaged spin potential and spin-shear stress (where the averaging is performed only over fluid cells with energy density greater than $\varepsilon_{\rm sw}$) for three different interactions in the fireball of Au+Au collisions at 200 GeV, with an impact parameter of $b=8.4$ fm corresponding to 20-50% centrality [54]. At late times, the relaxation of the spin potential towards thermal vorticity is governed by the longest timescale, which is $\tau_{\omega}$ for all three interactions. At early times, the evolution is primarily governed by the transport coefficients $\tau_{\kappa}$, $\mathfrak{d}$, $\tau_{\mathfrak{t}}$ and $\ell_{\kappa\mathfrak{t}}$. For $I_{S}$ and $I_{SP}$ cases, smaller values of $\tau_{\kappa}$ make the relaxation term dominant, whereas the peak value of the rise is controlled by $\mathfrak{d}$ and $\tau_{\mathfrak{t}}$. However, for the $I_{V}$ case, we observe significant overdamping at early times due to relatively large values of $\tau_{\kappa}$, $\tau_{\mathfrak{t}}$, and $\mathfrak{d}$ compared to the other two interactions. To rule out any numerical artifacts, we performed a series of numerical experiments by tuning individual transport coefficients, and conclude that this is indeed a physical effect. The relatively large value of $\tau_{\kappa}$ reduces the relaxation term’s contribution compared to the term involving the coefficient $\ell_{\kappa\mathfrak{t}}$, which has a damping effect. This damping persists until the spin-shear stress $\mathfrak{t}^{\mu\nu}$ nearly vanishes at later times, allowing the relaxation term to dominate and drive the system by increasing toward the thermal vorticity.
@@
With a qualitative understanding of spin evolution, we now apply our framework to compute the spin polarization of $\Lambda$ hyperons. We first examine the individual contributions of $\omega^{\mu}_{0}$, $\kappa^{\mu}_{0}$, and $\mathfrak{t}^{\mu\nu}$ to $\Lambda$ polarization (see Fig. 8 in Appendix F ) for the three different interactions. We find that the transverse momentum dependence of the $y$-component of polarization exhibits marginal sensitivity to the interaction type, with the spin-shear stress contribution being small at low $p_{T}$ but increasing at higher $p_{T}$. In contrast, for the azimuthal angle dependence of the $z$-component, the contribution from $\omega_{0}^{\mu}$ is small, and that from $\kappa_{0}^{\mu}$ is largely independent of the interaction type. Notably, the spin-shear stress plays a significant role, exhibiting strong sensitivity to the interaction type and behaving similar to the thermal shear in Ref. [14].
@@
The net local $\Lambda$ polarization, obtained by summing the individual contributions from $\omega^{\mu}_{0}$, $\kappa^{\mu}_{0}$, and $\mathfrak{t}^{\mu\nu}$, for the interactions considered in this work, is shown in Fig. 2. The results are compared to experimental data (teal dots). There is rather good agreement with the experimental data on the pseudorapidity ($\eta$) and the transverse momentum ($p_{T}$) dependence of the $y$-component of the polarization for the $I_{SP}$ interaction. Our approach also correctly reproduces the sign of the longitudinal polarization for all three interactions. Regarding the utility of spin hydrodynamics in the context of HIC, these results are very encouraging. Our work considers a fully dynamical system initialized with zero spin current, which is subsequently generated and transported through the exchange between orbital and spin angular momentum. In the course of this process, we obtain the correct sign of longitudinal polarization, with the tensor $\mathfrak{t}^{\mu\nu}$ in spin hydrodynamics playing a role analogous to the thermal shear ($\xi^{\mu\nu}$) in local-equilibrium scenarios. For the collision energy considered in this work ($\sqrt{s_{NN}}=200$ GeV), we conclude that the assumption of the spin thermalizing to its equilibrium value (given through the thermal vorticity) by the end of the fireball’s lifetime is a reasonable approximation, in agreement with the conclusion of Ref. [24]. However, at lower energies, this may no longer be the case. Crucially, the framework of spin hydrodynamics treated in this work should be applicable to these energies as well, since it captures the non-equilibrium spin dynamics.

### v2

@@ Abstract
We present a framework for spin dynamics in the quark-gluon plasma created in relativistic heavy-ion collisions. Under the approximation of small polarization, macroscopic spin degrees of freedom decouple from the background, and their evolution equations and transport coefficients have been computed using quantum kinetic theory of massive particles with nonlocal collisions. Employing this theory, we numerically solve dissipative relativistic spin hydrodynamics. We explore three interaction scenarios between constituent particles and apply this framework to compute both global and local spin polarization of $\Lambda$ hyperons in Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV. Our results show that the initially vanishing spin potential relaxes toward thermal vorticity, driving global polarization. Furthermore, we demonstrate that the sign of longitudinal polarization is sensitive to the interaction type, emphasizing the need for a consistent treatment of dissipative effects in spin hydrodynamics to describe experimental data.
@@
I Introduction
The study of spin polarization in relativistic heavy-ion collisions (HICs) has gained considerable interest in recent years, owing to the observation of global and local polarization of hyperons [1, 2, 3, 4, 5, 6, 7, 8]. Spin polarization observables provide probes of the vorticity and spin dynamics in the hot and dense fireball formed in relativistic HIC. The earliest attempt using a perturbative QCD-inspired model [9] to understand global polarization attributed it to the large orbital angular momentum, which polarizes the quarks in the same direction through spin-orbit coupling, analogous to the well-known Barnett effect [10].
In the quantum-statistical approach, the spin degrees of freedom are assumed to be in local thermodynamic equilibrium at the fireball freezeout, making it possible to define a density operator and derive an analytical expression for spin polarization at leading order in terms of the thermal vorticity, $\varpi^{\mu\nu}=\frac{1}{2}(\partial^{\nu}\beta^{\mu}-\partial^{\mu}\beta^{\nu})$, where $\beta^{\mu}=u^{\mu}/T$ is the inverse four-temperature, with $u^{\mu}$ denoting the fluid four-velocity and $T$ the temperature [11]. The spin polarization generated solely from thermal vorticity shows very good agreement with experimental data on the global polarization of $\Lambda$-hyperons across Beam Energy Scan collision energies [12]. However, the data on local polarization, particularly the longitudinal polarization, exhibits an opposite sign compared to hydrodynamic predictions. It was soon realized that the thermal shear, defined as $\xi^{\mu\nu}=\frac{1}{2}(\partial^{\nu}\beta^{\mu}+\partial^{\mu}\beta^{\nu})$, provides an additional contribution to spin polarization [13, 14, 15, 16, 17, 18, 19], with a sign opposite to that of thermal vorticity. The combined effect of thermal vorticity and thermal shear leads to a net negative longitudinal polarization. However, the correct sign and magnitude of the observed $\Lambda$ polarization in Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV can be reproduced only under the isothermal approximation, where gradients of temperature on the freezeout surface are neglected [14].
@@ I Introduction
While the equations of spin hydrodynamics can be formulated in several ways [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56], the most popular approach is to derive them from quantum kinetic theory (QKT), which describes a quantum-field theoretical system in the limit where the fields are sufficiently localized to be treated in terms of quasi-particles. Utilizing this framework, one possible approach consists in taking the spin tensor to be conserved [28, 29, 57, 58], implying that the energy-momentum tensor is symmetric. A numerical solution of spin hydrodynamics derived in such a way was presented in Ref. [59] with the spin tensor taken to be in de Groot, van Leeuwen, and van Weert (GLW) pseudogauge. For a suitable choice of initial spin potential, their results show that using a delayed initial time for spin hydrodynamics can successfully describe polarization measurements. This is suggestive of strong dissipative effects during the initial stages of the fireball that may be responsible for the relaxation of the spin potential toward its equilibrium value given by the thermal vorticity. Indeed, it is also possible to derive these dissipative effects in a quantum kinetic framework. In Ref. [26], it has been found that the relaxation of the spin potential toward the thermal vorticity is sourced by so-called nonlocal collisions [60, 61]. Intuitively, this is sensible since, in order to exchange orbital angular momentum and spin, the collision cannot occur at a single space-time point [29]. On a more technical side, these collisions essentially depend on the spin-changing matrix elements of the underlying theory, and the equations solved in Ref. [59] refer to a system in which these interactions can be neglected. In this work, we go beyond this approximation and numerically solve the equations of dissipative spin hydrodynamics. The hydrodynamic and spin degrees of freedom are evolved simultaneously at each timestep. Since our aim is to apply this framework to heavy-ion collisions, we work in the small polarization limit, where the spin sector does not significantly back-react on the hydrodynamic background, effectively decoupling the two sectors at leading order. The aim of this work is not to provide a full statistical analysis, but rather to demonstrate how successfully and consistently our framework can describe the spin-polarization data under the same hydrodynamic background conditions as those used in the equilibrium approach. This direct comparison allows us to assess the impact of going beyond the isothermal approximation and including dissipative spin dynamics.
@@
II Theoretical background
The degrees of freedom quantifying the spin density of the fluid can be described via the spin potential $\Omega^{\mu\nu}_{0}$, which constitutes an antisymmetric rank-two tensor and can thus be decomposed into electric and magnetic parts, [^1] Note that the sign of $\kappa_{0}^{\mu}$ differs between Ref. [29] and this work, which uses the same conventions as Refs. [26, 55].
@@ II Theoretical background
Furthermore, when considering dissipative spin hydrodynamics, it has been shown in Ref. [55] that one has to consider an additional rank-2 tensor $\mathfrak{t}^{\mu\nu}$ which has the same symmetries as the shear-stress tensor $\pi^{\mu\nu}$ and can thus be referred to as the spin-shear stress. In the GLW pseudogauge, it appears in the spin tensor as $S^{\mu,\alpha\beta}\sim\mathfrak{t}^{\mu[\alpha}u^{\beta]}$. [^2] Note that the equations of motion for the spin degrees of freedom $\omega_{0}^{\mu}$, $\kappa_{0}^{\mu}$, and $\mathfrak{t}^{\mu\nu}$ follow from (a truncation of) the moments of the quantum-kinetic Boltzmann equation and are independent of the definition of macroscopic currents such as the spin tensor.
@@
Figure 1: Time evolution of the energy-weighted space-averaged (left) xzxz component of the spin potential, and (right) zzzz component of the spin-shear stress, for three different interactions. Due to the symmetry of our initial condition model, the spatial average of all other components of the spin potential is zero.
The relaxation times $\tau_{\omega}$ and $\tau_{\kappa}$ quantify the time scales on which $\omega_{0}^{\mu}$ and $\kappa_{0}^{\mu}$ approach their asymptotic values, i.e., the first terms on the right-hand sides of Eqs. (2a) and (2b). As mentioned in the introduction, these times are (inversely) proportional to the nonlocal collision terms. On the other hand, the coefficient $\tau_{\mathfrak{t}}$, which sets the time scale of $\mathfrak{t}^{\mu\nu}$, is (inversely) proportional to local collisions, indicating that the relaxation of the spin-shear stress is not driven solely through spin-orbit coupling. The asymptotic value of the spin-shear stress, given by the first term on the right-hand side of Eq. (2c), is proportional to the shear tensor, reminiscent of the thermal-shear contribution in Ref. [13]. The corresponding transport coefficient $\mathfrak{d}$ is given by a ratio of nonlocal and local collision terms. In particular, while it depends on the underlying microscopic interaction, it is independent of its absolute strength, and vanishes when the collisions are local.
III Numerical framework
@@
Solving Eqs. (2a)-(2c) requires $T$, $u^{\mu}$, and their gradients from the background hydrodynamics. The details of the background hydrodynamics are discussed in Appendix A. We use the code developed in Refs. [59, 63, 64] to solve the equations of background hydrodynamics numerically in (3+1)-dimensions in Milne coordinates $(\tau,x,y,\eta_{s})$, where $\tau=\sqrt{t^{2}-z^{2}}$ and $\eta_{s}=\frac{1}{2}\ln\left(\frac{t+z}{t-z}\right)$. The code, developed in Forran, incorporates the Godunov-type relativistic Harten-Lax-van Leer-Einfeldt (HLLE) approximate Riemann solver for the computation of numerical fluxes at fluid cell boundaries, as described in Refs. [65, 66]. The implementation of Eqs. (2a)-(2c) is inspired by the publicly available vHLLE code [66, 67]. Additionally, the FORTRAN code supports parallel programming on multiple threads using OpenMP [68], a feature not present in the public version of vHLLE. The equations of background hydrodynamics are closed with an equation of state for which we use the lattice-QCD-based equation of state NEOS-BQS [69, 70, 71, 72, 73, 74]. We extend the code to solve the spin equations, Eqs. (2a)-(2c), with the numerical details provided in Appendix D.
@@
For the computation of the transport coefficients in Eqs. (2a)-(2c), one has to provide an explicit model of the system. Deferring the use of a kinetic framework with all particle species included to future work, we employ a simplified treatment by assuming a system of spin-$\nicefrac{{1}}{{2}}$ particles with a mass of 300 MeV, which is on the order of the thermal masses $m_{\text{th}}\sim gT$ of the particles in the hot and dense fireball. We consider three types of interactions among the particles, namely (i) scalar, where the interaction Lagrangian is $\mathcal{L}_{\text{int,S}}=G(\bar{\psi}\psi)^{2}$, (ii) scalar+pseudoscalar, where $\mathcal{L}_{\text{int,SP}}=G[(\bar{\psi}\psi)^{2}-(\bar{\psi}\gamma_{5}\psi)^{2}]$, [^3] This interaction is equivalent to the well-known Nambu-Jona–Lasinio (NJL) model [75]. and (iii) vector, with $\mathcal{L}_{\text{int,V}}=-G(\bar{\psi}\gamma^{\mu}\psi)^{2}$, with a coupling constant $G$. We denote these three scenarios by $I_{S}$, $I_{SP}$, and $I_{V}$, respectively. As mentioned earlier, the transport coefficient $\mathfrak{d}$ depends only on the form of the interaction but is independent of the coupling constant $G$. The same can be said for the relaxation times $\tau_{\omega}$, $\tau_{\kappa}$, and $\tau_{\mathfrak{t}}$, if we express them in units of the shear-stress relaxation time $\tau_{\pi}$. Consequently, we tabulate the relaxation times $\tau_{\omega}$, $\tau_{\kappa}$, and $\tau_{\mathfrak{t}}$ as functions of $z=m/T$ in units of $\tau_{\pi}$ for the three types of interactions mentioned above (see Fig. 4). All other transport coefficients in the spin equations are tabulated as functions of $z$ in units of the corresponding spin relaxation times and isotropic pressure to make them dimensionless (see Ref. [55] for details). We assume that these dimensionless ratios are universal, depending only on the nature of the interaction between particles, and thus remain unchanged for the quark-gluon fluid formed in relativistic HIC. Consequently, for transport coefficients that involve pressure, we use the local thermodynamic equilibrium pressure, $P$, of the fireball to extract their numerical values from tables of dimensionless ratios. This approach aligns with standard hydrodynamic simulations, such as MUSIC [76], where second-order transport coefficients computed using kinetic theory in the massless limit are assumed to apply to the quark-gluon fluid. To solve Eqs. (2a)-(2c), we initialize $\omega_{0}^{\mu}$, $\kappa_{0}^{\mu}$, and $\mathfrak{t}^{\mu\nu}$ to zero. This choice of initial condition for the spin degrees of freedom assumes that the total angular momentum at the beginning of the hydrodynamic evolution of the fireball is dominated solely by the orbital component.
@@
Figure 2: Comparison of results from dissipative spin hydrodynamics with experimental data for three interaction scenarios for Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV. The simulations are performed with fixed impact parameter, $b=8.4$ fm, corresponding to 20-50% centrality. The top panel shows the (left) pseudorapidity and (right) transverse momentum dependence of the $y$ component of $\Lambda$ polarization. The bottom panel shows the (left) azimuthal and (right) transverse momentum dependence of the $z$ component of $\Lambda$ polarization. Experimental data are taken from Refs. [2, 3] with the updated decay parameter ($\alpha_{\Lambda}=0.732$).
@@
The spin vector in the rest frame of $\Lambda$, $S^{*\mu}=(0,\textbf{S}^{*})$ is obtained from $S^{\mu}=(S^{0},\textbf{S})$ as follows:
@@ IV Results
With the framework described above, we analyze the behavior of dissipative spin hydrodynamics in HICs. In Fig. 1, we show the time evolution of the energy-weighted, space-averaged spin potential and spin-shear stress (where the averaging is performed only over fluid cells with energy density greater than $\varepsilon_{\rm sw}$) for three different interactions in the fireball of Au+Au collisions at 200 GeV, with an impact parameter of $b=8.4$ fm corresponding to 20-50% centrality [59]. At late times, the relaxation of the spin potential toward thermal vorticity is governed by the longest timescale, which is $\tau_{\omega}$ for all three interactions. At early times, the evolution is primarily governed by the transport coefficients $\tau_{\kappa}$, $\mathfrak{d}$, $\tau_{\mathfrak{t}}$ and $\ell_{\kappa\mathfrak{t}}$. For $I_{S}$ and $I_{SP}$ cases, smaller values of $\tau_{\kappa}$ make the relaxation term dominant, whereas the peak value of the rise is controlled by $\mathfrak{d}$ and $\tau_{\mathfrak{t}}$. However, for the $I_{V}$ case, we observe significant overdamping at early times due to relatively large values of $\tau_{\kappa}$, $\tau_{\mathfrak{t}}$, and $\mathfrak{d}$ compared with the other two interactions. To rule out any numerical artifacts, we performed a series of numerical experiments by tuning individual transport coefficients, and concluded that this is indeed a physical effect. The relatively large value of $\tau_{\kappa}$ reduces the relaxation term’s contribution compared with the term involving the coefficient $\ell_{\kappa\mathfrak{t}}$, which has a damping effect. This damping persists until the spin-shear stress $\mathfrak{t}^{\mu\nu}$ nearly vanishes at later times, allowing the relaxation term to dominate and drive the system by increasing toward the thermal vorticity.
@@
With a qualitative understanding of spin evolution, we now apply our framework to compute the spin polarization of $\Lambda$ hyperons. We first examine the individual contributions of $\omega^{\mu}_{0}$, $\kappa^{\mu}_{0}$, and $\mathfrak{t}^{\mu\nu}$ to $\Lambda$ polarization (see Fig. 8) for the three different interactions. We find that the transverse momentum dependence of the $y$ component of polarization exhibits marginal sensitivity to the interaction type, with the spin-shear stress contribution being small at low $p_{T}$ but increasing at higher $p_{T}$. In contrast, for the azimuthal angle dependence of the $z$ component, the contribution from $\omega_{0}^{\mu}$ is small, and that from $\kappa_{0}^{\mu}$ is largely independent of the interaction type. Notably, the spin-shear stress plays a significant role, exhibiting strong sensitivity to the interaction type and behaving similar to the thermal shear in Ref. [14].
@@
The net local $\Lambda$ polarization, obtained by summing the individual contributions from $\omega^{\mu}_{0}$, $\kappa^{\mu}_{0}$, and $\mathfrak{t}^{\mu\nu}$, for the interactions considered in this work, is shown in Fig. 2. The results are compared with experimental data (teal dots). There is rather good agreement with the experimental data on the pseudorapidity ($\eta$) and the transverse momentum ($p_{T}$) dependence of the $y$ component of the polarization for the $I_{SP}$ interaction. It is to be noted that our current implementation includes only (quasi) particles and not anti-(quasi) particles. As a result, a discrepancy arises between our results and the experimental data on $p_{T}$ dependence of $P_{z}$, which corresponds to a combined average of $\Lambda+\bar{\Lambda}$. Our approach also correctly reproduces the sign of the longitudinal polarization for all three interactions. Regarding the utility of spin hydrodynamics in the context of HIC, these results are very encouraging. Our work considers a fully dynamical system initialized with zero spin current, which is subsequently generated and transported through the exchange between orbital and spin angular momentum. In the course of this process, we obtain the correct sign of longitudinal polarization, with the tensor $\mathfrak{t}^{\mu\nu}$ in spin hydrodynamics playing a role analogous to the thermal shear ($\xi^{\mu\nu}$) in local-equilibrium scenarios. For the collision energy considered in this work ($\sqrt{s_{NN}}=200$ GeV), we conclude that the assumption of the spin thermalizing to its equilibrium value (given through the thermal vorticity) by the end of the fireball’s lifetime is a reasonable approximation, in agreement with the conclusion of Ref. [26]. However, at lower energies, this may no longer be the case. Crucially, the framework of spin hydrodynamics treated in this work should be applicable to these energies as well, since it captures the non-equilibrium spin dynamics.
V Summary